_____ 5. What statistical test would you use to analyze the data derived from this study?

A common belief held by laypersons is that smoking marijuana affects a person’s pupil size. To study this issue, Weil, Zinberg, and Nelson (1968) administered a high dose of marijuana to ten men by having them smoke a potent marijuana cigarette. These men were between the ages of 21 and 26 and all smoked tobacco cigarettes regularly, but had never smoked marijuana. Pupil size was measured under constant illumination before and after each man smoked the marijuana cigarette. Measurements were taken in millimeters. The means and variance information is reported below. Based on these data, can one conclude that marijuana increases pupil size?

A. One sample t test

B. Independent samples t test

C. Correlated samples t test

D. One-Way Between Subjects ANOVA

_____ 6. When sample sizes are small, power will tend to be _____, everything else being equal.

A. low

B. moderate

C. high

D. there is no effect

_____ 7. What statistical test would you use to analyze the data derived from this study?

An organizational psychologist is investigating the effects of shift type on job satisfaction. There are three shift types: standard (five 8-hour days, two days off), concentrated (four 10-hour days, three days off), and extended (10 consecutive 8-hour days followed by four days off). In this study, twenty-four workers are randomly assigned to one of the three conditions and job satisfaction is measured after six months.

A. One sample t test

B. Independent samples t test

C. Correlated samples t test

D. One-Way Between Subjects ANOVA

_____ 8. If a Between Subjects ANOVA study involves 4 groups and a total of 24 subjects, the df between is ____ and the df within is ___.

A. 3; 18

B. 2:20

C. 3; 20

D. 2; 24
Answer #1

5.

Since the pupil size of same people were measured before and after smoking the marijuana cigarette, this is an

C. Correlated samples t test

6.

When sample sizes are small, standard error tend to be high and the critical region to reject the null hypothesis would decrease. Thus, Power (Reject H0 when null hypothesis is false) will tend to decrease.

A. low

7.

Since the factor levels of treatments (shift type) on job satisfaction is more than two, we will use One-Way Between Subjects ANOVA.

D. One-Way Between Subjects ANOVA

8.

df Between = Number of groups - 1 = 4 - 1 = 3

df within = Number of subjects - Number of groups = 24 - 4 = 20

C. 3; 20